How many electrons make up a charge of -35.0 uC?
Answer #1

Charge of a single electron (q) = -1.6 x 10-19 C

Given charge (Q) = -35 \mu C = -35 x 10-6 C

Let n be the number of electrons, then:

nq = Q

n = Q/q = (-35 x 10-6/-1.6 x 10-19 ) = 2.1875 x 1014 electrons